if constant a in damped motion is given by if constant a in damped motion is given by 10. what is the value of time constant10. what is the value of time constant
1
Expert's answer
2020-11-27T14:02:27-0500

A damped motion is described by


"x(t)=A_0e^{-\frac{b}{2m}t}."

The time constant is


"\tau=\frac{2m}{b},"

where "m" - mass of the bob,

"b" - damping constant.

Thus, the time constant is


"\tau=\frac{2m}{10}=\frac{m}{5}."